# Monochromatic products in random integer sets

### v1

@@ 1. Introduction
A collection of (ordered) subsets $\mathcal{S}\subset{\mathbb{N}}$ is partition-regular [^1] This definition is often stated in an infinitary fashion, asking for any $r$-colouring of $\mathbb{N}$} to give monochromatic $S\in\mathcal{S}$. When every member $S\in\mathcal{S}$ is finite, these two formulations are equivalent by the so-called compactness principle. if for any $r\in{\mathbb{N}}$, any $r$-colouring of $n]:=\{1,2\ldots,n\}$ with $n\in{\mathbb{N}}$ is sufficiently large, results in a monochromatic set $S\in\mathcal{S}$. Determining which arithmetic configurations lead to collections of sets $\mathcal{S}$ that are partition-regular is a major theme in modern combinatorics. The earliest result in this area of arithmetic Ramsey theory (and indeed in Ramsey theory as a whole) is due to Schur [SchurUpperBound] over a century ago, who showed that the collection of all sums $\{(a,b,c)\in{\mathbb{N}}^{3}:a+b=c\}$ is partition-regular. In 1933, Rado [rado1933studien] proved a far-reaching generalisation of Schur’s result, characterising for which matrices $A$ the system of linear homogeneous equations $A\mathbf{x}=0$ has a solution set that is partition-regular. This celebrated result in particular implies both Schur’s theorem and the famous theorem of van der Waerden [van1927beweis] showing that for any $r$, the collection of $r$-term arithmetic progressions $\{(a,a+d,\ldots,a+(r-1)d):a,d\in{\mathbb{N}}\}$ is partition-regular. Further work completely settled partition-regularity for collections of sets defined by finite sets of equations, see e.g. the survey of Hindman [Hindman_survey].
@@ 1.1. Random sets of integers
In the 1990s, researchers started to strengthen statements showing partition-regularity by exploring the Ramsey properties of sparse host sets. In the context of monochromatic sums, we define a set $A\subseteq{\mathbb{N}}$ to be $r$-Schur if any $r$-colouring of $A$ results in monochromatic $(a,b,c)\in A^{3}$ with $a+b=c$. When $r=2$, we simply refer to the set $A$ as being Schur. Whilst Schur’s theorem [SchurUpperBound] guarantees that $[n]$ is Schur for all $n$ sufficiently large, it turns out that much sparser subsets of $[n]$ are typically also Schur. This was shown in pioneering work of Graham, Rödl and Ruciński [SumThreshold]. The set $n]_{p}$ denotes the binomial random subset of $[n]$ obtained by keeping each integer in $[n]$ with probability $p=p(n)$ independently and we say that a property $\mathcal{A}\subseteq 2^{[n]}$ holds in $n]_{p}$ asymptotically almost surely (a.a.s. for short) if the probability that it holds tends to 1 as $n$ tends to infinity.
Theorem 1.1 ([SumThreshold]).
@@
For any property $\mathcal{A}\subseteq 2^{[n]}$, a function $p_{*}=p_{*}(n)$ is the [^2] This is a slight abuse of notation as thresholds are not unique but rather determined up to constant factors. threshold for $\mathcal{A}$ if whenever $p=\omega(p_{*})$, we have that $n]_{p}$ a.a.s. has the property $\mathcal{A}$ whilst whenever $p=o(p_{*})$, we have that $n]_{p}$ a.a.s. does not have the property $\mathcal{A}$. Thus Theorem 1.1 establishes that $n^{-1/2}$ is the threshold for the Schur property. Further work established a sparse random analogue of Rado’s theorem [rado1933studien], determining the threshold for $n]_{p}$ to contain monochromatic solutions to sets of homogeneous equations $A{\bf{x}}=0$ in any colouring, for any $A$ such that the solution set is partition-regular. The $(0)$-statements for these thresholds as well as $(1)$-statements for a subclass of so-called density-regular matrices $A$ were proven by Rödl and Ruciński [RadoPartition] who conjectured that their $(0)$-statements provided the correct threshold in all cases. This took over a decade to verify with Friedgut, Rödl and Schacht [friedgut2010ramsey] finally providing a proof for all $(1)$-statements. The $(1)$-statements for all these Rado-type properties in fact hold for any bounded number of colours $r$ and, perhaps surprisingly, the thresholds do not depend on the number of colours $r$. In particular, it follows from [friedgut2010ramsey] that for any $r\geq 2$, the threshold for $n]_{p}$ to be $r$-Schur is at $n^{1/2}$.
@@
Shortly after the breakthrough work of Friedgut, Rödl and Schacht [friedgut2010ramsey] providing $(1)$-statements for the random Rado theorem, Conlon and Gowers [2016.CG] and independently Schacht [2016.Schacht] developed powerful transference principles which establish optimal $(1)$-statements for a wide range of extremal and Ramsey properties in random discrete structures. In particular, these provided new proofs for some of $(1)$-statements for Rado-type problems discussed above. By adopting an abstract viewpoint of independent sets in hypergraphs, Balogh, Morris and Samotij [2015.BMS] and independently Saxton and Thomason [2015.ST] then developed the theory of hypergraph containers which encapsulated many of the applications of the previous transference principles and led to an extremely flexible and elegant tool for establishing $(1)$-statements in random discrete structures. In particular, Nenadov and Steger [nenadov2016short] showed that one can use hypergraph containers to establish thresholds for Ramsey properties. This has led to a golden age in random Ramsey theory with many recent advances. Restricting only to the arithmetic setting here, we mention thresholds for solutions with repeated entries to Rado-type equations [spiegel2017note], resilience versions of Rado thresholds [hancock2019independent], perturbed Schur thresholds [aigner2019monochromatic, das2024schur], asymmetric random Rado theorems [aigner2023asymmetric, hancock2022asymmetric], sharp thresholds for Schur and van der Waerden [friedgut2015sharp, friedgut2022sharp, schulenburg2016threshold], thresholds for Rado type properties in random subsets of the primes and abelian groups [freschi2024typical] and canonical van der Waerden thresholds [alvarado2024canonical].
@@ 1.2. Monochromatic products
All of the random arithmetic Ramsey results mentioned above concern solutions to linear equations. In this paper, we make a first foray into exploring non-linear equations by considering monochromatic products. The fact that the set [^3] We omit the integer 1 here to avoid trivialities. $\{(a,b,c)\in([n]\setminus\{1\})^{3}:ab=c\}$ is partition-regular can be derived from Schur’s theorem for sums by considering powers of 2. In general, we say a set $A\subseteq{\mathbb{N}}$ is $r$-product-Schur if any $r$-colouring of $S$ results in monochromatic $(a,b,c)\in A^{3}$ with $ab=c$ and when $r=2$, we will just say that $S$ is product-Schur. As the property of being product-Schur is monotone increasing [^4] That is, if $A$ is product-Schur and $A\subseteq B$ then $B$ is also product-Schur., the Bollobás-Thomason theorem [bollobas1987threshold] implies that there exists a threshold for the property. Our main theorem gives bounds on this threshold.
@@
Theorem 1.2 makes a first step towards answering a recent question of Mattos, Mergoni Cecchelli and Parczyk [mattos2025product]. Their motivation came from a question of Prendiville at the 2022 British Combinatorial Conference (BCC), which asked for another strengthening of Schur’s theorem to be generalised to the setting of products. Answering a $\mathdollar 100$ question of Graham [SumThreshold], Robertson and Zeilberger [robertson19982] and independently Schoen [schoen1999number] proved that any 2-colouring of $[n]$ gives $(\tfrac{1}{11}+o(1))n^{2}$ monochromatic sums and that this is tight (their results are listed with the constant $1/22$ as they consider $(a,b,c)$ and $(b,a,c)$ as the same solution). For more colours, determining asymptotically the minimum number of monochromatic sums is an intriguing open question. Prendiville asked how many monochromatic products are necessarily guaranteed when colouring $n]\setminus\{1\}$. This was then addressed by Aragão, Chapman, Ortega and Souza [aragao2023number] and independently by Mattos, Mergoni Cecchelli and Parczyk [mattos2025product]. For two colours, the first group [aragao2023number] determined the minimum number of products precisely, showing that it is $(\tfrac{1}{2\sqrt{2}}+o(1))n^{1/2}\log n$. The second group [mattos2025product] showed a lower bound of $n^{1/3-o(1)}$ for this problem and also initiated the study of other variations of Schur’s theorem in the product setting. In particular, they determined precisely the threshold for the appearance of a product in the random set $n]_{p}$ and asked the question of determining the threshold for the random set to be product-Schur.
@@
The collection of products in $n]\setminus\{1\}$ behaves very differently to the collection of sums. Indeed there are $\Theta(n\log n)$ products which is much less than the $\Theta(n^{2})$ sums in $[n]$, and the collection is less regular, with the number of products involving an integer $a\in[n]\setminus\{1\}$ depending heavily on the location of $a$ in the interval. In the Ramsey setting, the so-called Schur multiplicity results discussed above also show a difference in behaviour between the sum and product case. Indeed for sums when $n$ is large we have that a positive fraction of all sums are monochromatic for any 2-colouring of $[n]$, an example of a phenomenon known as supersaturation. For products on the other hand, there are colourings in which only a vanishing fraction of the $\Theta(n\log n)$ products are monochromatic. Given this, it is perhaps not surprising that the proof of our (1)-statement for Theorem 1.2 differs from existing proofs in random Ramsey theory. Indeed, the transference principles [2016.CG, 2016.Schacht] and the hypergraph container method [2015.BMS, 2015.ST] discussed above crucially rely on supersaturation and thus seem to be powerless in the setting of monochromatic products. To prove our (1)-statement, we instead appeal to a certain parameterised family of small configurations which have the product-Schur property and show that a.a.s. one of these configurations will appear in our random set $n]_{p}$. This is reminiscent of the proof of the $(1)$-statement for Schur’s theorem in the randomly perturbed model [aigner2019monochromatic, das2024schur] when starting with a dense set of integers. The configurations we use are also similar (in fact contain) those used in [aragao2023number] to count monochromatic solutions. Known proof methods establishing $(0)$-statements for random Ramsey results also seem ineffective in the product setting and we establish the (0)-statement of Theorem 1.2 through a careful greedy colouring procedure.
@@ 1.3. More colours
We also give initial bounds on thresholds for the random set to be $r$-product-Schur for $r\geq 3$. In particular, these bounds show that, unlike all of the thresholds [friedgut2010ramsey] for solutions to sets of linear equations as in Rado’s theorem, for monochromatic products, the threshold depends on the number of colours. To state our results, we introduce some definitions. The Schur number $S(r)$ is the minimum $n$ such that any $r$-colouring of $[n]$ results in a monochromatic sum. The double-sum Schur number $S^{\prime}(r)$ is the minimum $n$ such that any $r$-colouring of $[n]$ results in either a monochromatic sum $(a,b,c)\in[n]^{3}$ with $a+b=c$ or a monochromatic shifted sum $(a,b,c)\in[n]^{3}$ with a+b=c1a+b=c-1. Double-sum Schur numbers were introduced by Abbott and Hanson [AbbottHanson] and also feature in the work of Mattos, Mergoni Cecchelli and Parczyk [mattos2025product] bounding the size of the largest $r$-product-Schur subset of $[n]$.
@@
One can check by hand or by computer that $S^{\prime}(2)=S(2)=5$ and $S^{\prime}(3)=S(3)=15$. The bounds given by Theorem 1.4 are thus weaker than those from Theorem 1.2 for two colours, whilst for three colours Theorem 1.4 shows that the threshold for being $3$-product-Schur lies between $n^{-1/14}$ and $n^{-1/196}$, and in particular is larger than the threshold for being $2$-product Schur. In fact, for a small number of colours we can do slightly better than the bound given by Theorem 1.4 (0), showing that the threshold for being $3$-product Schur is at least $n^{-1/19}$, see Remark 4.2. In general, the values of $S^{\prime}(r)$ and $S(r)$ may differ, for example $S^{\prime}(4)=41<45=S(4)$ but we have that $3S(r-1)-1\leq S^{\prime}(r)\leq S(r)$ as shown by Abbott and Hanson [AbbottHanson, Theorem 4.1]. Calculating (double-sum) Schur numbers is extremely challenging already for $r=5$ [heule2018schur] and the best general bounds give that $\Omega(3.17176^{r})\leq S(r)\leq r!e$ with the lower bound due to Exoo [SchurLowerBound] and the upper bound already shown by Schur himself [SchurUpperBound]. Theorem 1.4 thus gives that if $p^{*}_{r}$ is the threshold for $n]_{p}$ being $r$-product Schur, then $n^{-1/c_{1}^{r}}\leq p_{r}^{*}\leq n^{-1/c_{2}^{r\log r}}$ for some constants $c_{1},c_{2}>0$.
@@ 2. The (0)-statement for two colours
The precise exponent of $D(n)$ here is not relevant, we just take it large enough to give us enough spare room in later bounds. The divisor bound tells us that $D(n)=n^{o(1)}$ (see, for example, Theorem 13.12 in [polylogbound]), so the (1)-statement is a consequence of the following.
@@
If $p\leq n^{-1/9}\omega(n)^{-1}$, then a.a.s. we have that $n]_{p}$ is product-Schur.

### v2

@@ 1. Introduction
A collection of (ordered) subsets $\mathcal{S}\subset{\mathbb{N}}$ is partition-regular [^1] This definition is often stated in an infinitary fashion, asking for any $r$-colouring of $\mathbb{N}$} to give monochromatic $S\in\mathcal{S}$. When every member $S\in\mathcal{S}$ is finite, these two formulations are equivalent by the so-called compactness principle. if for any $r\in{\mathbb{N}}$, any $r$-colouring of $n]:=\{1,2\ldots,n\}$ with $n\in{\mathbb{N}}$ is sufficiently large, results in a monochromatic set $S\in\mathcal{S}$. Determining which arithmetic configurations lead to collections of sets $\mathcal{S}$ that are partition-regular is a major theme in modern combinatorics. The earliest result in this area of arithmetic Ramsey theory (and indeed in Ramsey theory as a whole) is due to Schur [30] over a century ago, who showed that the collection of all sums $\{(a,b,c)\in{\mathbb{N}}^{3}:a+b=c\}$ is partition-regular. In 1933, Rado [23] proved a far-reaching generalisation of Schur’s result, characterising for which matrices $A$ the system of linear homogeneous equations $A\mathbf{x}=0$ has a solution set that is partition-regular. This celebrated result in particular implies both Schur’s theorem and the famous theorem of van der Waerden [32] showing that for any $r$, the collection of $r$-term arithmetic progressions $\{(a,a+d,\ldots,a+(r-1)d):a,d\in{\mathbb{N}}\}$ is partition-regular. Further work completely settled partition-regularity for collections of sets defined by finite sets of equations, see e.g. the survey of Hindman [20].
@@ 1.1. Random sets of integers
In the 1990s, researchers started to strengthen statements showing partition-regularity by exploring the Ramsey properties of sparse host sets. In the context of monochromatic sums, we define a set $A\subseteq{\mathbb{N}}$ to be $r$-Schur if any $r$-colouring of $A$ results in monochromatic $(a,b,c)\in A^{3}$ with $a+b=c$. When $r=2$, we simply refer to the set $A$ as being Schur. Whilst Schur’s theorem [30] guarantees that $[n]$ is Schur for all $n$ sufficiently large, it turns out that much sparser subsets of $[n]$ are typically also Schur. This was shown in pioneering work of Graham, Rödl and Ruciński [16]. The set $n]_{p}$ denotes the binomial random subset of $[n]$ obtained by keeping each integer in $[n]$ with probability $p=p(n)$ independently and we say that a property $\mathcal{A}\subseteq 2^{[n]}$ holds in $n]_{p}$ asymptotically almost surely (a.a.s. for short) if the probability that it holds tends to 1 as $n$ tends to infinity.
Theorem 1.1 ([16]).
@@
For any property $\mathcal{A}\subseteq 2^{[n]}$, a function $p_{*}=p_{*}(n)$ is the [^2] This is a slight abuse of notation as thresholds are not unique but rather determined up to constant factors. threshold for $\mathcal{A}$ if whenever $p=\omega(p_{*})$, we have that $n]_{p}$ a.a.s. has the property $\mathcal{A}$ whilst whenever $p=o(p_{*})$, we have that $n]_{p}$ a.a.s. does not have the property $\mathcal{A}$. Thus Theorem 1.1 establishes that $n^{-1/2}$ is the threshold for the Schur property. Further work established a sparse random analogue of Rado’s theorem [23], determining the threshold for $n]_{p}$ to contain monochromatic solutions to sets of homogeneous equations $A{\bf{x}}=0$ in any colouring, for any $A$ such that the solution set is partition-regular. The $(0)$-statements for these thresholds as well as $(1)$-statements for a subclass of so-called density-regular matrices $A$ were proven by Rödl and Ruciński [25] who conjectured that their $(0)$-statements provided the correct threshold in all cases. This took over a decade to verify with Friedgut, Rödl and Schacht [15] finally providing a proof for all $(1)$-statements. The $(1)$-statements for all these Rado-type properties in fact hold for any bounded number of colours $r$ and, perhaps surprisingly, the thresholds do not depend on the number of colours $r$. In particular, it follows from [15] that for any $r\geq 2$, the threshold for $n]_{p}$ to be $r$-Schur is at $n^{1/2}$.
@@
Shortly after the breakthrough work of Friedgut, Rödl and Schacht [15] providing $(1)$-statements for the random Rado theorem, Conlon and Gowers [9] and independently Schacht [27] developed powerful transference principles which establish optimal $(1)$-statements for a wide range of extremal and Ramsey properties in random discrete structures. In particular, these provided new proofs for some of $(1)$-statements for Rado-type problems discussed above. By adopting an abstract viewpoint of independent sets in hypergraphs, Balogh, Morris and Samotij [7] and independently Saxton and Thomason [26] then developed the theory of hypergraph containers which encapsulated many of the applications of the previous transference principles and led to an extremely flexible and elegant tool for establishing $(1)$-statements in random discrete structures. In particular, Nenadov and Steger [22] showed that one can use hypergraph containers to establish thresholds for Ramsey properties. This has led to a golden age in random Ramsey theory with many recent advances. Restricting only to the arithmetic setting here, we mention thresholds for solutions with repeated entries to Rado-type equations [31], resilience versions of Rado thresholds [17], perturbed Schur thresholds [2, 10], asymmetric random Rado theorems [3, 18], sharp thresholds for Schur and van der Waerden [13, 14, 29], thresholds for Rado type properties in random subsets of the primes and abelian groups [12] and canonical van der Waerden thresholds [4].
@@ 1.2. Monochromatic products
All of the random arithmetic Ramsey results mentioned above concern solutions to linear equations. In this paper, we make a first foray into exploring non-linear equations by considering monochromatic products. The fact that the set [^3] We omit the integer 1 here to avoid trivialities. $\{(a,b,c)\in([n]\setminus\{1\})^{3}:ab=c\}$ is partition-regular can be derived from Schur’s theorem for sums by considering powers of 2. In general, we say a set $A\subseteq{\mathbb{N}}$ is $r$-product-Schur if any $r$-colouring of $S$ results in monochromatic $(a,b,c)\in A^{3}$ with $ab=c$ and when $r=2$, we will just say that $S$ is product-Schur. As the property of being product-Schur is monotone increasing [^4] That is, if $A$ is product-Schur and $A\subseteq B$ then $B$ is also product-Schur., the Bollobás-Thomason theorem [8] implies that there exists a threshold for the property. Our main theorem gives bounds on this threshold.
@@
Theorem 1.2 makes a first step towards answering a recent question of Mattos, Mergoni Cecchelli and Parczyk [21]. Their motivation came from a question of Prendiville at the 2022 British Combinatorial Conference (BCC), which asked for another strengthening of Schur’s theorem to be generalised to the setting of products. Answering a $\mathdollar 100$ question of Graham [16], Robertson and Zeilberger [24] and independently Schoen [28] proved that any 2-colouring of $[n]$ gives $(\tfrac{1}{11}+o(1))n^{2}$ monochromatic sums and that this is tight (their results are listed with the constant $1/22$ as they consider $(a,b,c)$ and $(b,a,c)$ as the same solution). For more colours, determining asymptotically the minimum number of monochromatic sums is an intriguing open question. Prendiville asked how many monochromatic products are necessarily guaranteed when colouring $n]\setminus\{1\}$. This was then addressed by Aragão, Chapman, Ortega and Souza [6] and independently by Mattos, Mergoni Cecchelli and Parczyk [21]. For two colours, the first group [6] determined the minimum number of products precisely, showing that it is $(\tfrac{1}{2\sqrt{2}}+o(1))n^{1/2}\log n$. The second group [21] showed a lower bound of $n^{1/3-o(1)}$ for this problem and also initiated the study of other variations of Schur’s theorem in the product setting. In particular, they determined precisely the threshold for the appearance of a product in the random set $n]_{p}$ and asked the question of determining the threshold for the random set to be product-Schur.
@@
The collection of products in $n]\setminus\{1\}$ behaves very differently to the collection of sums. Indeed there are $\Theta(n\log n)$ products which is much less than the $\Theta(n^{2})$ sums in $[n]$, and the collection is less regular, with the number of products involving an integer $a\in[n]\setminus\{1\}$ depending heavily on the location of $a$ in the interval. In the Ramsey setting, the so-called Schur multiplicity results discussed above also show a difference in behaviour between the sum and product case. Indeed for sums when $n$ is large we have that a positive fraction of all sums are monochromatic for any 2-colouring of $[n]$, an example of a phenomenon known as supersaturation. For products on the other hand, there are colourings in which only a vanishing fraction of the $\Theta(n\log n)$ products are monochromatic. Given this, it is perhaps not surprising that the proof of our (1)-statement for Theorem 1.2 differs from existing proofs in random Ramsey theory. Indeed, the transference principles [9, 27] and the hypergraph container method [7, 26] discussed above crucially rely on supersaturation and thus seem to be powerless in the setting of monochromatic products. To prove our (1)-statement, we instead appeal to a certain parameterised family of small configurations which have the product-Schur property and show that a.a.s. one of these configurations will appear in our random set $n]_{p}$. This is reminiscent of the proof of the $(1)$-statement for Schur’s theorem in the randomly perturbed model [2, 10] when starting with a dense set of integers. The configurations we use are also similar (in fact contain) those used in [6] to count monochromatic solutions. Known proof methods establishing $(0)$-statements for random Ramsey results also seem ineffective in the product setting and we establish the (0)-statement of Theorem 1.2 through a careful greedy colouring procedure.
@@ 1.3. More colours
We also give initial bounds on thresholds for the random set to be $r$-product-Schur for $r\geq 3$. In particular, these bounds show that, unlike all of the thresholds [15] for solutions to sets of linear equations as in Rado’s theorem, for monochromatic products, the threshold depends on the number of colours. To state our results, we introduce some definitions. The Schur number $S(r)$ is the minimum $n$ such that any $r$-colouring of $[n]$ results in a monochromatic sum. The double-sum Schur number $S^{\prime}(r)$ is the minimum $n$ such that any $r$-colouring of $[n]$ results in either a monochromatic sum $(a,b,c)\in[n]^{3}$ with $a+b=c$ or a monochromatic shifted sum $(a,b,c)\in[n]^{3}$ with a+b=c1a+b=c-1. Double-sum Schur numbers were introduced by Abbott and Hanson [1] and also feature in the work of Mattos, Mergoni Cecchelli and Parczyk [21] bounding the size of the largest $r$-product-Schur subset of $[n]$.
@@
One can check by hand or by computer that $S^{\prime}(2)=S(2)=5$ and $S^{\prime}(3)=S(3)=15$. The bounds given by Theorem 1.4 are thus weaker than those from Theorem 1.2 for two colours, whilst for three colours Theorem 1.4 shows that the threshold for being $3$-product-Schur lies between $n^{-1/14}$ and $n^{-1/196}$, and in particular is larger than the threshold for being $2$-product Schur. In fact, for a small number of colours we can do slightly better than the bound given by Theorem 1.4 (0), showing that the threshold for being $3$-product Schur is at least $n^{-1/19}$, see Remark 4.2. In general, the values of $S^{\prime}(r)$ and $S(r)$ may differ, for example $S^{\prime}(4)=41<45=S(4)$ but we have that $3S(r-1)-1\leq S^{\prime}(r)\leq S(r)$ as shown by Abbott and Hanson [1, Theorem 4.1]. Calculating (double-sum) Schur numbers is extremely challenging already for $r=5$ [19] and the best general bounds give that $\Omega(3.17176^{r})\leq S(r)\leq r!e$ with the lower bound due to Exoo [11] and the upper bound already shown by Schur himself [30]. Theorem 1.4 thus gives that if $p^{*}_{r}$ is the threshold for $n]_{p}$ being $r$-product Schur, then $n^{-1/c_{1}^{r}}\leq p_{r}^{*}\leq n^{-1/c_{2}^{r\log r}}$ for some constants $c_{1},c_{2}>0$.
@@ 2. The (0)-statement for two colours
The precise exponent of $D(n)$ here is not relevant, we just take it large enough to give us enough spare room in later bounds. The divisor bound tells us that $D(n)=n^{o(1)}$ (see, for example, Theorem 13.12 in [5]), so the (1)-statement is a consequence of the following.
@@
If $p\leq n^{-1/9}\omega(n)^{-1}$, then a.a.s. we have that $n]_{p}$ is not product-Schur.